OpenStudy (jazzyluvsu123):

1/5x=3/10 i do Divide 3/10 / 1/5 ?? do i flip the 1/5 to 5/1 ??

OpenStudy (anonymous):

1/8 + x = 1/2 x = 1/2 - 1/8

OpenStudy (jazzyluvsu123):

and my answer would be 3/8 ??

OpenStudy (anonymous):

x = 4/8 - 1/8 x = 3/8

OpenStudy (anonymous):

yes

OpenStudy (beastboy17):

I believe that is correct

OpenStudy (jazzyluvsu123):

1/5x=3/10 i do Divide 3/10 / 1/5 ?? do i flip the 1/5 to 5/1 ??